A speedboat moving at 30.0 m/s approaches a ship 100 m ahead. The pilot slows the boat with a constant acceleration of -3.50 m/s2. How long does it take the boat to reach the ship?
1
Expert's answer
2020-10-07T10:14:55-0400

When the boat is moving with constant acceleration, its position depends on time as

"\displaystyle s = v_0t + \frac{at^2}{2}"

Let's put into it "s=100\, m", "v_0=30.0 \, m/s" , "a = -3.50 \,m/s^2".

"100 = 30t -1.75 t^2"

"1.75 t^2 - 30t + 100=0"

"\displaystyle t_{1,2} = \frac{30 \pm \sqrt{900-700}}{3.5}= \frac{30\pm14.14}{3.5}=4.53 \;s"

We have chosen minus sign here because of the properties of parabola. s(t) is parabola. We start at some point, move to the apex of a parabola and then return back. We need to chose the first time when t satisfies the equation, namely, the smaller root.

Answer: "t = 4.53\,s"